OpenStudy (anonymous):

Change of variables double integral problem: double integral over region R (x+y)e^(x^2-y^2) dA with R being a rectangle enclosed by the lines...

OpenStudy (anonymous):

x-y=0, x-y=2, x+y=0, and x+y=3......

OpenStudy (anonymous):

So far I have u=x^2-y^2 and v=x+y, and solved change of variables via x=1/2(root u) +v and y=1/2(v-u) with double integral ve^u, but now I'm stuck on finding the boundaries and am wondering if I am doing it right?